# Radial and vertical constraints on the icy origin of HCO in the HD 163296 Protoplanetary Disk

### v1

@@ 3.2 Spatial Distribution of HCO Emission
3.3 Radial Distribution of HCO Emission
To have a more precise characterization of the spatial distribution of the emission, we deproject the zeroth-moment maps using the radial_profile function implemented in GoFish. The resulting deprojected and azimuthally averaged radial intensity profiles are shown in Figure 3, where the continuum profiles are in logarithmic scales to identify the outer edge of the millimeter continuum () more clearly. We defined as the radius where the emission falls below . Considering the two continuum deprojected radial profiles, we estimated an value of au, consistent with previous results at higher angular resolution (Carney et al., 2017; Law et al., 2021a). On the other hand, the radial profiles of the emission lines are generally more extended than those of the continuum, reaching distances up to au in the case of the four brightest transitions (i.e., , , , and ).
In terms of morphology, most of the profiles replicate the more prominent radial substructures already witnessed in previous high angular resolution studies of HCO in HD 163296 (e.g., Pegues et al., 2020; Guzmán et al., 2021): a central depression followed by a bright peak at au that then decreases with radius but undergoes two less prominent ring-like structures located at au (spatially coincident with ) and au. Regarding continuum morphology, a well-known set of rings and gaps has been previously characterized in the innermost au by higher angular resolution maps (e.g., Huang et al., 2018; Law et al., 2021a), which, unfortunately, are not resolved by our observations.
3.4 HCO Excitation Conditions and OPR
Since we have multiple ortho and para transitions that cover a wide range of upper-state energies ( K), we model simultaneously the emission of each line to determine the excitation temperature, the total (ortho + para) column density, and the OPR of HCO. Using the Planck function , the line intensity can be estimated from the excitation temperature, , and the line optical depth, , as follows:
@@ 3.4 HCO Excitation Conditions and OPR
HCO Line Frequency (GHz) (s) (K) (cm) ALMA band 3 (o) ALMA band 6 (p) (o) ALMA band 7 (p) (p) (p) (o)
Note. — The rest frequency and Einstein coefficient are taken from the JPL database (Pickett et al., 1998), while the upper-state energy, the upper-state degeneracy and the collisional coefficients to compute the critical densities are extracted from the Leiden Atomic and Molecular Database (LAMDA; Schöier et al., 2005).
@@
where is the rest frequency of the line, is the Einstein coefficient, is the velocity line width define as the FWHM of the line profile, and is the column density of the molecule in the upper-state of the transition. If we assume that the line emission occurs under local thermodynamic equilibrium (LTE), then
where is the total column density, is the partition function of HCO, and with are the energy and the degeneracy associated to the upper-state level of each transition, respectively (see Table 3). This should be a reasonable assumption for HD 163296 since the critical densities of the observed transitions ( cm, see Table 3) are lower than the typical gas densities estimated by models (Qi et al., 2011; Zhang et al., 2021) at those vertical heights where we expect to find HCO (e.g., Paneque-Carreño et al., 2023). In that sense, the excitation temperature can be used as a direct tracer of the kinetic gas temperature in those disk layers where HCO is emitting.
We defined separate partition functions for the ortho and para isomers. Using the same convention adopted by Terwisscha van Scheltinga et al. (2021), we defined:
where and are the upper-state degeneracy and energy of a state. The sum in Equations (4) and (5) is performed over the rotational ground states of a specific isomer (ortho or para) taken from the ExoMol database (Al-Refaie et al., 2015; Wang et al., 2020) but using state degeneracies consistent with the ones from the Leiden and Molecular Database (LAMDA; Schöier et al., 2005) since the former assumes an OPR of (see Terwisscha van Scheltinga et al., 2021, for more details).
By putting Equations (2) and (3) into Equation (1), and assuming a Gaussian line profile centered at km s (Teague et al., 2019) with a certain , we can model the line intensity as a function of and . As we have different partition functions, and is not necessarily equal to , we could in principle do a separate line modeling for each isomer and then obtain the ortho-to-para ratio . However, in this work we assume the same excitation temperature since we detected only two o-HCO lines. Hence, we fitted simultaneously the detected ortho and para transitions using the following relations:
where is the total column density of HCO considering both isomers. Thus, the free parameters to fit are , , , and . To fit the excitation conditions and the line width, we construct a likelihood function from Equation (1) and use the emcee package (Foreman-Mackey et al., 2013) to explore the parameter space and get the posterior probability distributions.
@@ 3.4.1 Disk-averaged Analysis
We first model the line emission averaged over the full radial extent of HCO ( au) in order to derive the disk-averaged excitation conditions. To account for the velocity structure of the disk, we used the average_spectrum function from GoFish, which corrects the spectrum for Keplerian rotation before averaging by using the stellar mass and the distance to the disk. For simplicity, we assume that the emitting layer is in the midplane (), which should be reasonable based on the estimates of the vertical height of HCO in HD 163296 (see Section 3.5 for further details). The retrieved disk-averaged spectra are shown in Figure 4, where the best-fit Gaussian models are overlayed with different colors for each HCO isomer. The best-fit excitation conditions and their corresponding uncertainties are deduced from the 16th, 50th and 84th percentiles of the posterior distributions, whose results are cm, K, and for a Gaussian line width of km s. The disk-averaged line optical depths are between , verifying an optically thin regime. We also checked that these excitation conditions are consistent with the non-detection of the line, as its predicted line intensity is lower than the sensitivity of the observations.
@@
As shown in Figure 4, the LTE predictions underestimate the disk-averaged emission from the line wings, indicating that the line profiles are probably more Lorentzian than pure Gaussian. Consequently, we performed the same analysis but using a Voigt profile. We find an improvement in the fit quality for the line wings, and recover slightly higher values for and the OPR, although consistent with the results from the Gaussian fitting when considering the uncertainties (see Appendix B). A similar line profile behavior was found in TW Hya using CO line observations, that were attributed to pressure broadening (Yoshida et al., 2022). Our HCO results may hint at something similar, but further analysis is needed to investigate the origin of the non-Gaussian profiles in HD 163296.
@@
Previous efforts have been made to constrain the distribution of HCO across HD 163296. For the disk-integrated column density, Carney et al. (2019) estimated a value of about cm, which is lower by a factor of than our derived value. However, it is still consistent, considering that they assumed a higher excitation temperature of K. Moreover, to compute the total column density, they used integrated flux measurements of a single line from SMA observations with lower angular resolution (Qi et al., 2013). Thus, the differences between the disk-averaged estimates could be due to choosing a different area of integration, or filtering out more extended emission (either in ALMA or the SMA).
@@
Regarding the excitation temperature, Carney et al. (2017) estimated a lower limit of K on the kinetic temperature using three p-HCO rotational transitions. Nevertheless, as they did not have higher energy transitions (e.g., ), and their two weakest lines were detected only through a matched filter analysis (see Loomis et al., 2018b, for details of this technique), the temperature estimation was uncertain. Later, a similar result was found by Guzmán et al. (2018), where a rotational temperature of K was obtained using three o-HCO transitions with lower angular resolution SMA data, consistent with our results considering the uncertainties.
@@
In the case of the ortho-to-para ratio, the only known determination for HD 163296 was made also by Guzmán et al. (2018), who complemented the three o-HCO transitions mentioned above with three p-HCO transitions and estimated an ortho-to-para ratio between , depending on the vertical height of the emission. This range of values agrees with our OPR determination. However, since the derived uncertainties are purely statistical and do not consider the assumptions mentioned in Section 3.4, we cannot rule out a disk-averaged OPR value consistent with in HD 163296 (see Section 4.2.3).
@@ 3.4.2 Radially Resolved Analysis
Since we can resolve the variation of the HCO emission with the disk radius (see Figure 3), we repeat the same line fitting procedure but using the deprojected radial spectra instead of the disk-averaged one to derive the excitation conditions as a function of radius. Using GoFish, we applied the radial_spectra function to deproject the spectra into different radial bins. Similarly to the disk-averaged case, the emission is shifted and stacked according to the Keplerian motion. However, it is now azimuthally averaged over different radial annuli whose width corresponds to of the beam major axis. In this case, we construct a likelihood function for each radial bin to determine the best-fit models at different radii, as shown in Figure 5.
Figure 6 shows the resulting deprojected radial profiles of the excitation conditions and the disk-averaged values for comparison, whereas radial profiles are shown in Appendix C. In general, the mean values of each profile are compatible with the disk-averaged results, and we confirmed that the emission is optically thin throughout the disk (see Figure 9). The derived profile behaves similar to those from previous studies where the emission of HCO was resolved in HD 163296 (Pegues et al., 2020; Guzmán et al., 2021). Total column density values range from to cm within a radius of au, displaying two remarkable features: a central gap followed by a global peak close to the CO snowline located at au (Qi et al., 2015, corrected by Gaia distance) and a secondary ring or plateau near au, as defined by our continuum observations.
Regarding the excitation temperature, this is the first time it has been resolved for HCO in HD 163296 since, previously, it was computed through a disk-averaged analysis (Carney et al., 2017; Guzmán et al., 2018) or assumed uniform across the disk (Pegues et al., 2020; Guzmán et al., 2021) either due to a lack of sufficient angular resolution or multiple line observations. We find that the profile decreases smoothly from to K, where a temperature of K is reached close to CO snowline and K at .
On the other hand, the ortho-to-para ratio profile is relatively constant across the disk, mostly centered at but with uncertainties that remain compatible with the statistically expected value of . Nevertheless, the profile exhibits small modulations that are apparently consistent with some of the HCO radial substructures witnessed by Law et al. (2021a) that are spatially resolved by our observations, which could be linked with changes in the chemistry.
@@ 3.5 Vertical Height of the Emission
Although the radial distribution of the excitation conditions gives us valuable clues about the formation of HCO in protoplanetary disks, the latter have a three-dimensional structure where the vertical distribution of the emission is also an essential factor to have a complete understanding of the contribution of the different formation pathways. To complement our radially resolved analysis, we used the Python package ALFAHOR666https://github.com/teresapaz/alfahor to constrain empirically the vertical height of HCO in HD 163296 from the observations (Paneque-Carreño et al., 2023).
By having the stellar and geometrical properties of the disk, ALFAHOR computes a vertical profile based on a geometrical method (Pinte et al., 2018) that relies on the location of the emission maxima inside a user-defined mask for each channel (see Paneque-Carreño et al., 2023, for more details). Using the stellar and disk parameters from Table 1, and masking manually the channel emission for each HCO transition, we derived the vertical profiles shown in Figure 7. Since this empirical method requires an appropriate combination of angular resolution and signal-to-noise ratio to distinguish the near and far “wings” of the channel emission (see Figure 1 in Paneque-Carreño et al., 2023), we applied this analysis only to the four brightest transitions in our sample.
HCO is apparently tracing an emission surface close to . Still, as the lines used in this work have , we cannot discard the presence of HCO at higher disk layers since optically thin emission makes it challenging to isolate surfaces due to back- and front-side confusion across the disk. Thus, the determined value should be interpreted as a lower limit. We did not notice remarkable differences in the height predicted from different lines, which agrees with the LTE assumption where all transitions have the same excitation temperature since the disk gas temperature changes with the radius and the height. Similar to the integrated intensities (see Figure 3), the vertical profiles present radial modulations, as previously reported by Paneque-Carreño et al. (2023), where the lowest heights are spatially coincident with the HCO rings. This suggests that the structures seen from the radial profiles are likely related to changes in the HCO chemistry, which could be related to the slight radial variations of the OPR (see Section 4.2 for further discussion).
@@ 4 Discussion
4.1 The chemistry of HCO in Protoplanetary Disks
The reliability of HCO as a tracer of the icy O reservoir in protoplanetary disks will depend on its dominant formation pathway since, at typical disk conditions, it could be formed by gas-phase and grain-surface chemistry (Walsh et al., 2014). According to laboratory experiments (e.g., Fockenberg & Preses, 2002; Atkinson et al., 2006) and chemical modeling (e.g., van der Marel et al., 2014; Loomis et al., 2015), HCO is formed efficiently in the gas-phase at low temperatures ( K) by the following neutral-neutral reaction:
@@ 4.1 The chemistry of HCO in Protoplanetary Disks
On the other hand, in those cold regions where most of the O-containing carriers are frozen, HCO and other more complex species (e.g., CHOH) are expected to form by the successive hydrogenation of CO on dust grains (e.g., Watanabe & Kouchi, 2002; Fuchs et al., 2009):
@@
which should be relevant in the vicinity and inside the regions with abundant CO ices, delimited by the CO snowline and the CO snow surface (Öberg et al., 2017; Guzmán et al., 2021). Since thermal desorption of HCO occurs at warmer conditions than the CO freeze-out temperature (Pegues et al., 2020; Terwisscha van Scheltinga et al., 2021), non-thermal processes are required to desorb the grain-surface HCO into the gas-phase.
@@
Based on the above, the key takeaway from previous works is that probably a combination of gas-phase and grain-surface formation pathways is needed to explain the emission morphologies deduced from HCO observations (e.g., Pegues et al., 2020; Guzmán et al., 2021), which is supported by predictions from chemical models that include both kinds of chemistry (Loomis et al., 2015; Öberg et al., 2017) but also by empirical measurements of vertical stratification in disks (Podio et al., 2020; Paneque-Carreño et al., 2023). Still, most previous studies lack multiple transitions and different isomer observations, and therefore, the excitation conditions with the OPR, which are directly related to HCO formation, are barely constrained.
@@
In fact, there are only two disks where the above quantities are now well-resolved. Terwisscha van Scheltinga et al. (2021) recently resolved the excitation conditions and the OPR of HCO in the protoplanetary disk around TW Hya. They found that HCO emits from a relatively warm ( K) layer, and a decreasing OPR with radius, going from values consistent with 3 within 60 au (which corresponds to the extent of the pebble disk) and values of 2 beyond this radius. The relatively uniform excitation conditions and OPR gradient suggests a single gas-phase formation pathway dominates the formation of HCO across the TW Hya disk, although a contribution from grain-surface chemistry cannot be completely ruled-out (Terwisscha van Scheltinga et al., 2021). The other case is HD 163296, with the observations presented in this work.
4.2 Constraints on HCO Formation in HD 163296
4.2.1 Radial Distribution of HCO

### v2

@@ 3.2 Spatial Distribution of H2CO Emission
3.3 Radial Distribution of H2CO Emission
To have a more precise characterization of the spatial distribution of the emission, we deproject the zeroth-moment maps using the radial_profile function implemented in GoFish. The resulting deprojected and azimuthally averaged radial intensity profiles are shown in Figure 3, where the continuum profiles are in logarithmic scales to identify the outer edge of the millimeter continuum () more clearly. We defined as the radius where the emission falls below . Considering the two continuum-deprojected radial profiles, we estimated an value of au, consistent with previous results at higher angular resolution (Carney et al., 2017; Law et al., 2021a). On the other hand, the radial profiles of the emission lines are generally more extended than those of the continuum, reaching distances up to au in the case of the four brightest transitions (i.e., , , , and ).
In terms of morphology, most of the profiles replicate the more prominent radial substructures already witnessed in previous high angular resolution studies of H2CO in HD 163296 (e.g., Pegues et al., 2020; Guzmán et al., 2021): a central depression followed by a bright peak at au that then decreases with radius but undergoes two less prominent ring-like structures located at au (spatially coincident with ) and au. Regarding continuum morphology, a well-known set of rings and gaps has been previously characterized in the innermost au by higher angular resolution maps (e.g., Huang et al., 2018; Law et al., 2021a), which, unfortunately, are not resolved by our observations.
3.4 H2CO Excitation Conditions and OPR
Since we have multiple ortho and para transitions that cover a wide range of upper-state energies ( K), we simultaneously model the emission of each line to determine the excitation temperature, the total (ortho + para) column density, and the OPR of H2CO. Using the Planck function , the line intensity can be estimated from the excitation temperature, , and the line optical depth, , as follows:
@@ 3.4 H2CO Excitation Conditions and OPR
H2CO Line Frequency (GHz) (s-1) (K) (cm-3) ALMA Band 3 (o) ALMA Band 6 (p) (o) ALMA Band 7 (p) (p) (p) (o)
Note. — The rest frequency and Einstein coefficient are taken from the JPL database (Pickett et al., 1998), while the upper-state energy, the upper-state degeneracy, and the collisional coefficients to compute the critical densities are extracted from the Leiden Atomic and Molecular Database (LAMDA; Schöier et al., 2005).
@@
where is the rest frequency of the line, is the Einstein coefficient, is the velocity line width define as the FWHM of the line profile, and is the column density of the molecule in the upper state of the transition. If we assume that the line emission occurs under local thermodynamic equilibrium (LTE), then
where is the total column density, is the partition function of H2CO, and with are the energy and the degeneracy associated to the upper-state level of each transition, respectively (see Table 3). This should be a reasonable assumption for HD 163296 since the critical densities of the observed transitions ( cm-3, see Table 3) are lower than the typical gas densities estimated by models (Qi et al., 2011; Zhang et al., 2021) at those vertical heights where we expect to find H2CO (e.g., Paneque-Carreño et al., 2023). In that sense, the excitation temperature can be used as a direct tracer of the kinetic gas temperature in those disk layers where H2CO is emitting.
We defined separate partition functions for the ortho and para isomers. Using the same convention adopted by Terwisscha van Scheltinga et al. (2021), we defined
where and are the upper-state degeneracy and energy of a state. The sum in Equations (4) and (5) is performed over the rotational ground states of a specific isomer (ortho or para) taken from the ExoMol database (Al-Refaie et al., 2015; Wang et al., 2020) but using state degeneracies consistent with the ones from LAMDA (Schöier et al., 2005) since the former assumes an OPR of (see Terwisscha van Scheltinga et al., 2021, for more details).
By putting Equations (2) and (3) into Equation (1), and assuming a Gaussian line profile centered at km s-1 (Teague et al., 2019) with a certain , we can model the line intensity as a function of and . As we have different partition functions, and is not necessarily equal to , we could in principle do a separate line modeling for each isomer and then obtain the ortho-to-para ratio . However, in this work, we assume the same excitation temperature since we detected only two o-H2CO lines. Hence, we fitted simultaneously the detected ortho and para transitions using the following relations:
where is the total column density of H2CO considering both isomers. Thus, the free parameters to fit are , , , and . To fit the excitation conditions and the line width, we construct a likelihood function from Equation (1) and use the emcee package (Foreman-Mackey et al., 2013) to explore the parameter space and get the posterior probability distributions.
@@ 3.4.1 Disk-averaged Analysis
We first model the line emission averaged over the full radial extent of H2CO ( au) in order to derive the disk-averaged excitation conditions. To account for the velocity structure of the disk, we used the average_spectrum function from GoFish, which corrects the spectrum for Keplerian rotation before averaging by using the stellar mass and the distance to the disk. For simplicity, we assume that the emitting layer is in the midplane (), which should be reasonable based on the estimates of the vertical height of H2CO in HD 163296 (see Section 3.5 for further details). The retrieved disk-averaged spectra are shown in Figure 4, where the best-fit Gaussian models are overlaid with different colors for each H2CO isomer. The best-fit excitation conditions and their corresponding uncertainties are deduced from the 16th, 50th, and 84th percentiles of the posterior distributions, whose results are cm-2, K, and for a Gaussian line width of km s-1. The disk-averaged line optical depths are between and , verifying an optically thin regime. We also checked that these excitation conditions are consistent with the nondetection of the line, as its predicted line intensity is lower than the sensitivity of the observations.
@@
As shown in Figure 4, the LTE predictions underestimate the disk-averaged emission from the line wings, indicating that the line profiles are probably more Lorentzian than pure Gaussian. Consequently, we performed the same analysis but using a Voigt profile. We find an improvement in the fit quality for the line wings, and recover slightly higher values for and the OPR, although consistent with the results from the Gaussian fitting when considering the uncertainties (see Appendix B). A similar line profile behavior was found in TW Hya using 12CO line observations, that were attributed to pressure broadening (Yoshida et al., 2022). Our H2CO results may hint at something similar, but further analysis is needed to investigate the origin of the non-Gaussian profiles in HD 163296.
@@
Previous efforts have been made to constrain the distribution of H2CO across HD 163296. For the disk-integrated column density, Carney et al. (2019) estimated a value of about cm-2, which is lower by a factor of than our derived value. However, it is still consistent, considering that they assumed a higher excitation temperature of K. Moreover, to compute the total column density, they used integrated flux measurements of a single line from Submillimeter Array (SMA) observations with lower angular resolution (Qi et al., 2013). Thus, the differences between the disk-averaged estimates could be due to choosing a different area of integration, or filtering out more extended emission (either in ALMA or the SMA).
@@
Regarding the excitation temperature, Carney et al. (2017) estimated a lower limit of K on the kinetic temperature using three p-H2CO rotational transitions. Nevertheless, as they did not have higher energy transitions (e.g., ), and their two weakest lines were detected only through a matched filter analysis (see Loomis et al., 2018b, for details of this technique), the temperature estimation was uncertain. Later, a similar result was found by Guzmán et al. (2018), where a rotational temperature of K was obtained using three o-H2CO transitions with lower angular resolution SMA data, consistent with our results considering the uncertainties.
@@
In the case of the ortho-to-para ratio, the only known determination for HD 163296 was also made by Guzmán et al. (2018), who complemented the three o-H2CO transitions mentioned above with three p-H2CO transitions and estimated an ortho-to-para ratio between and , depending on the vertical height of the emission. This range of values agrees with our OPR determination. However, since the derived uncertainties are purely statistical and do not consider the assumptions mentioned in Section 3.4, we cannot rule out a disk-averaged OPR value consistent with in HD 163296 (see Section 4.2.3).
@@ 3.4.2 Radially Resolved Analysis
Since we can resolve the variation of the H2CO emission with the disk radius (see Figure 3), we repeat the same line fitting procedure but use the deprojected radial spectra instead of the disk-averaged one to derive the excitation conditions as a function of radius. Using GoFish, we applied the radial_spectra function to deproject the spectra into different radial bins. Similarly to the disk-averaged case, the emission is shifted and stacked according to the Keplerian motion. However, it is now azimuthally averaged over different radial annuli whose width corresponds to one-quarter of the beam’s major axis. In this case, we construct a likelihood function for each radial bin to determine the best-fit models at different radii, as shown in Figure 5.
Figure 6 shows the resulting deprojected radial profiles of the excitation conditions and the disk-averaged values for comparison, whereas radial profiles are shown in Appendix C. In general, the mean values of each profile are compatible with the disk-averaged results, and we confirmed that the emission is optically thin throughout the disk (see Figure 9). The derived profile behaves similarly to those from previous studies where the emission of H2CO was resolved in HD 163296 (Pegues et al., 2020; Guzmán et al., 2021). Total column density values range from cm-2 within a radius of au, displaying two remarkable features: a central gap followed by a global peak close to the CO snowline located at au (Qi et al., 2015, corrected by Gaia distance) and a secondary ring or plateau near au, as defined by our continuum observations.
Regarding the excitation temperature, this is the first time it has been resolved for H2CO in HD 163296 since, previously, it was computed through a disk-averaged analysis (Carney et al., 2017; Guzmán et al., 2018) or assumed uniform across the disk (Pegues et al., 2020; Guzmán et al., 2021) either due to a lack of sufficient angular resolution or multiple-line observations. We find that the profile decreases smoothly from to K, where a temperature of K is reached close to CO snowline and K at .
On the other hand, the ortho-to-para ratio profile is relatively constant across the disk, mostly centered at but with uncertainties that remain compatible with the statistically expected value of . Nevertheless, the profile exhibits small modulations that are apparently consistent with some of the H2CO radial substructures witnessed by Law et al. (2021a) that are spatially resolved by our observations, which could be linked with changes in the chemistry.
@@ 3.5 Vertical Height of the Emission
Although the radial distribution of the excitation conditions gives us valuable clues about the formation of H2CO in protoplanetary disks, the latter has a three-dimensional structure where the vertical distribution of the emission is also an essential factor to having a complete understanding of the contribution of the different formation pathways. To complement our radially resolved analysis, we used the Python package ALFAHOR666https://github.com/teresapaz/alfahor to constrain empirically the vertical height of H2CO in HD 163296 from the observations (Paneque-Carreño et al., 2023).
By having the stellar and geometrical properties of the disk, ALFAHOR computes a vertical profile based on a geometrical method (Pinte et al., 2018) that relies on the location of the emission maxima inside a user-defined mask for each channel (see Paneque-Carreño et al., 2023, for more details). Using the stellar and disk parameters from Table 1, and masking manually the channel emission for each H2CO transition, we derived the vertical profiles shown in Figure 7. Since this empirical method requires an appropriate combination of angular resolution and signal-to-noise ratio to distinguish the near and far wings of the channel emission (see Figure 1 in Paneque-Carreño et al., 2023), we applied this analysis only to the four brightest transitions in our sample.
H2CO is apparently tracing an emission surface close to . Still, as the lines used in this work have , we cannot discard the presence of H2CO at higher disk layers since optically thin emission makes it challenging to isolate surfaces due to back- and front-side confusion across the disk. Thus, the determined value should be interpreted as a lower limit. We did not notice remarkable differences in the height predicted from different lines, which agrees with the LTE assumption where all transitions have the same excitation temperature since the disk gas temperature changes with the radius and the height. Similar to the integrated intensities (see Figure 3), the vertical profiles present radial modulations, as previously reported by Paneque-Carreño et al. (2023), where the lowest heights are spatially coincident with the H2CO rings. This suggests that the structures seen from the radial profiles are likely related to changes in the H2CO chemistry, which could be related to the slight radial variations of the OPR (see Section 4.2 for further discussion).
@@ 4 Discussion
4.1 The chemistry of H2CO in Protoplanetary Disks
The reliability of H2CO as a tracer of the icy O reservoir in protoplanetary disks will depend on its dominant formation pathway since, at typical disk conditions, it could be formed by gas-phase and grain-surface chemistry (Walsh et al., 2014). According to laboratory experiments (e.g., Fockenberg & Preses, 2002; Atkinson et al., 2006) and chemical modeling (e.g., van der Marel et al., 2014; Loomis et al., 2015), H2CO is formed efficiently in the gas phase at low temperatures ( K) by the following neutral-neutral reaction:
@@ 4.1 The chemistry of H2CO in Protoplanetary Disks
On the other hand, in those cold regions where most of the O-containing carriers are frozen, H2CO and other more complex species (e.g., CH3OH) are expected to form by the successive hydrogenation of CO on dust grains (e.g., Watanabe & Kouchi, 2002; Fuchs et al., 2009):
@@
which should be relevant in the vicinity and inside the regions with abundant CO ices, delimited by the CO snowline and the CO snow surface (Öberg et al., 2017; Guzmán et al., 2021). Since thermal desorption of H2CO occurs at warmer conditions than the CO freeze-out temperature (Pegues et al., 2020; Terwisscha van Scheltinga et al., 2021), nonthermal processes are required to desorb the grain-surface H2CO into the gas-phase.
@@
Based on the above, the key takeaway from previous works is that probably a combination of gas-phase and grain-surface formation pathways is needed to explain the emission morphologies deduced from H2CO observations (e.g., Pegues et al., 2020; Guzmán et al., 2021), which is supported by predictions from chemical models that include both kinds of chemistry (Loomis et al., 2015; Öberg et al., 2017) but also by empirical measurements of vertical stratification in disks (Podio et al., 2020; Paneque-Carreño et al., 2023). Still, most previous studies lack multiple transitions and different isomer observations, and therefore, the excitation conditions with the OPR, which are directly related to H2CO formation, are barely constrained.
@@
In fact, there are only two disks where the above quantities are now well-resolved. Terwisscha van Scheltinga et al. (2021) recently resolved the excitation conditions and the OPR of H2CO in the protoplanetary disk around TW Hya. They found that H2CO emits from a relatively warm ( K) layer, and a decreasing OPR with radius, going from values consistent with 3 within 60 au (which corresponds to the extent of the pebble disk) and values of 2 beyond this radius. The relatively uniform excitation conditions and OPR gradient suggest a single gas-phase formation pathway dominates the formation of H2CO across the TW Hya disk, although a contribution from grain-surface chemistry cannot be completely ruled out (Terwisscha van Scheltinga et al., 2021). The other case is HD 163296, with the observations presented in this work.
4.2 Constraints on H2CO Formation in HD 163296
4.2.1 Radial Distribution of H2CO